OpenStudy (anonymous):

I NEED HELP!! An energy bar company manufactures peanut butter and chocolate bars. Peanut butter bars cost 25¢ per lb. The chocolate bars cost 30¢ per lb. The company sold 260 lbs of bars for $71. How many of the bars sold were peanut butter?

OpenStudy (imstuck):

You need a system of equations to figure this out. You need to relate the number of bars sold to each other as the first equation, then you would need to relate the money they cost and how much money they made as the second equation.

OpenStudy (anonymous):

i got this equation.....25x+(260-x)30=71

OpenStudy (imstuck):

Call pounds of pb bars p; call pounds of chocolate bars c. pounds of pb bars + pounds of chocolate bars = 260 pounds. That's the first equation.

OpenStudy (imstuck):

p+c=260

OpenStudy (imstuck):

Next relate the money in a second equation.

OpenStudy (anonymous):

so 25p+(260-p)30c=71

OpenStudy (imstuck):

pb costs .25 per pound and c costs .30 per pound and there is a total of $71 made. Convert either the cents to dollars or the dollars to cents so you are working in only one denomination of money.

OpenStudy (imstuck):

We now have that .25pb + .30c = 71

OpenStudy (imstuck):

That's our second equation.

OpenStudy (imstuck):

p + c = 260 .25p + .30c = 71

OpenStudy (anonymous):

.25pb+.30c=71 is second and were trying to find the first

OpenStudy (imstuck):

Solve them simultaneously. You could use substitution for c; p = 260-c.

OpenStudy (anonymous):

yeah but how are we going to find the first equation

OpenStudy (imstuck):

Ok, let's do this step by step. We related the numbers of pounds of bars in our first equation, which was p + c = 260

OpenStudy (imstuck):

Our second equation relating the cost per pound and the money they made is .25p + .30c = 71

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right

OpenStudy (imstuck):

We need to solve these at the same time, using one to solve the other. If we use substitution we can eliminate one variable. From the first equation, p + c = 260 we can solve for p to get rid of c like this: p = 260 - c

OpenStudy (imstuck):

Now we have a "new value" for p that we can sub into our second equation to get that equation into terms of only p and no c. Like this:

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

.25p + .30c = 71 .25(260-c) + .30c = 71 I meant "no p only c" When we do this we are eliminating the problem of 2 variables. We get rid of the p and only need to worry about c.

OpenStudy (imstuck):

p = 260 - c, so we sub 260 - c in for p.

OpenStudy (anonymous):

(260)p+c=

OpenStudy (imstuck):

.25(260-c) + .30c = 71 .25 * 260 = 65....25 * -c = -.25c 65 - .25c + .30c = 71 is what we have now.

OpenStudy (imstuck):

Not just 260, it's "260 - c" that is equal to p. When we see the p in our second equation we need to replace it with 260-c. No more p then.

OpenStudy (anonymous):

gotchu

OpenStudy (imstuck):

65 - .25c + .30c = 71 We can now solve that for c. Can you do that?

OpenStudy (imstuck):

Combine the c terms on one side of the equals sign and the numbers on the other side of the equals sign.

OpenStudy (anonymous):

.55c

OpenStudy (anonymous):

65-.55c=71

OpenStudy (imstuck):

it's not .30c + .25c, it's .30c - .25c

OpenStudy (anonymous):

then it would be 0.05

OpenStudy (imstuck):

.30c - .25c = .05c You added the .25c and it's supposed to be subtracted.

OpenStudy (imstuck):

yes....05c. Go on...

OpenStudy (imstuck):

You now have .05c = 71-65 which is...

OpenStudy (imstuck):

.05c = ?

OpenStudy (anonymous):

.05c=6

OpenStudy (imstuck):

Good! Now solve for c. What do you get?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

unless its a frction, in that case it would be 6/.05

OpenStudy (imstuck):

Good! Verry good! That tells us that there was 120 pounds of chocolate bars sold. The total amount in pounds of chocolate and pb is 160. If we know that 120 pounds of chocolate was sold and the total pounds of pb + c = 160, and c = 120, then pb + 120 = 160. How do you solve for pb now?

OpenStudy (anonymous):

wouldnt it be 40?

OpenStudy (anonymous):

160-120=40

OpenStudy (imstuck):

good...120 chocolate and 40 pb. that's your answer!

OpenStudy (anonymous):

thank you so much for the help